OpenStudy (anonymous):

How to find this series:

OpenStudy (anonymous):

series from 1 to infinity of 1/((n+1)(n+2))

OpenStudy (michele_laino):

hint: we can write: \[\Large \frac{1}{{\left( {n + 1} \right)\left( {n + 2} \right)}} = \frac{1}{{n + 1}} - \frac{1}{{n + 2}}\]

OpenStudy (michele_laino):

so, we have: \[ \begin{gathered} \sum\limits_{n = 1}^m {\frac{1}{{\left( {n + 1} \right)\left( {n + 2} \right)}}} = \sum\limits_{n = 1}^m {\left( {\frac{1}{{n + 1}} - \frac{1}{{n + 2}}} \right)} = \hfill \\ \hfill \\ = \left( {\frac{1}{2} - \frac{1}{3}} \right) + \left( {\frac{1}{3} - \frac{1}{4}} \right) + \left( {\frac{1}{4} - \frac{1}{5}} \right) + ... - \frac{1}{{m + 1}} + \frac{1}{{m + 1}} - \frac{1}{{m + 2}} = \hfill \\ \hfill \\ = \frac{1}{2} - \frac{1}{{m + 2}} \hfill \\ \end{gathered} \] now what is the limit value of that sum when as m goes to +infinity?

OpenStudy (anonymous):

Telescoping series ftw! the limit approaches 1/2! Awesome ty